OpenStudy (anonymous):

Jennifer has been saving for college for 57 months. The first month, she saved $11. She was able to save more money each month than the month before. She ended up saving $19,779.00. How much more did she save each month? How would I solve this? Can someone please walk me through this?

OpenStudy (danjs):

have you seen or studied a geometric series of numbers, each term in the series is some constant value times the last tterm. I assume in this prob, the "how much more saved each month" is that constant value. the general form for this is y(x) = A * b^x value y = starting value A * (base) ^ (x months)

OpenStudy (danjs):

they give everything but the constant increase each month, the base of the exponent 19779 = 11 * (b)^57

OpenStudy (anonymous):

Well, I was absent to my class for the last few days and just received my makeup work. I don't believe this is a geometric series, but I may be wrong. The formula I have is [(av1+avn)/2](n)=total sum From that I tried... [11+(11+56(x))]/2 _____ = 19,779 but I couldn't figure out if that was right so far or if so what I should plug into the space. I tried attaching a picture, but I'm not sure if it went through. Sorry for my late reply, I'm on my phone.